cos(π/3 + A) = sin(π/6 - A)


1
Expert's answer
2021-03-04T13:49:59-0500

"\cos(\frac{\pi}{3} + A) = \cos\frac{\pi}{3} \cos A - \sin \frac{\pi}{3} \sin A = \sin\frac{\pi}{6} \cos A - \cos \frac{\pi}{6} \sin A = \sin (\frac{\pi}{6} - A)" as desired.